OpenStudy (anonymous):

A ball is thrown into the air with an upward velocity of 28 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 28t + 7. How long does it take the ball to reach its maximum height? What is the ball’s maximum height? Round to the nearest hundredth, if necessary 1.75 s; 7 ft 0.88 s; 43.75 ft 0.88 s; 17.5 ft 0.88 s; 19.25 ft

OpenStudy (across):

\[f(t)=-16t^2+28t+7\]\[f'(t)=-32t+28\implies -32t+28=0\implies t=\frac{7}{8}\]\[f(7/8)=19.25\]

OpenStudy (anonymous):

thank youu(: